OpenStudy (anonymous):

Find the perimeter of the following shape, rounded to the nearest tenth: Shape ABCD is shown. Point A is at negative 3, 5. Point B is at 1, 7. Point C is at 3, 3. Point D is at negative 1, 1. 9.8 16 17.9 24

OpenStudy (nirmalnema):

do u have any fig

OpenStudy (anonymous):

yea i do @nirmalnema here

OpenStudy (anonymous):

Okay

OpenStudy (nirmalnema):

A=(-3,5) B=1, 7. C=3, 3 D=-1, 1. NOW BY DISTANCE FORMULA U CAN FIND LENGTH OF SIDE AB , BC ,CD ,DA DISTANCE FORMULA= \[\sqrt{(x1−x2)^{2}+(y1−y2)^{2}}\]

OpenStudy (nirmalnema):

lengths are AB= \[\sqrt{20}\] BC= \[\sqrt{20}\] CD= \[\sqrt{20}\] DA= \[\sqrt{20}\]

OpenStudy (anonymous):

\[\sqrt{20}=4.47\] right

OpenStudy (nirmalnema):

NOW PERIMETER = SUM OF ALL THE SIDE LENGTH =\[\sqrt{20}+\sqrt{20}+ \sqrt{20}+ \sqrt{20}\] =\[4\sqrt{20}\] =4*4.47

OpenStudy (nirmalnema):

YUP U R RIGHT

OpenStudy (nirmalnema):

17.88

OpenStudy (anonymous):

the answer is C right because there is no 17.88

OpenStudy (nirmalnema):

yup.

OpenStudy (anonymous):

oh okay Thank YOU

OpenStudy (nirmalnema):

welcome..!!